# DegDiT: Controllable Audio Generation with Dynamic Event Graph Guided Diffusion Transformer

### v1

@@ III-B Training Strategy and Dataset Construction
$\mathbf{E}_{\text{cond}}=\textrm{Concat}[\mathbf{E}_{\text{text}},\mathbf{H},\mathbf{E}_{\text{dur}}].$ (9)
@@
$\mathbf{x}_{t}=(1-t)\mathbf{x}_{1}+t\mathbf{x}_{0},$ (10)
@@
$\mathbf{u}_{t}=\frac{d\mathbf{x}_{t}}{dt}=\mathbf{x}_{0}-\mathbf{x}_{1}.$ (11)
Enabling the model to learn a time-dependent vector field that transports samples from the noise distribution to the target audio distribution, we train the model $\epsilon_{\theta}(\mathbf{x}_{t},t,\mathbf{E}_{\text{cond}})$ to predict this velocity using the flow matching loss:
$\mathcal{L}_{\text{FM}}=\mathbb{E}_{\mathbf{x}_{1},\mathbf{x}_{0},t}\left[\|\epsilon_{\theta}(\mathbf{x}_{t},t,\mathbf{E}_{\text{cond}})-\mathbf{u}_{t}\|_{2}^{2}\right].$ (12)
@@
Let $N_{\mathcal{D}^{\prime}}$ denotes the predefined sample numbers of $\mathcal{D}^{\prime}$, with $0.25N_{\mathcal{D}^{\prime}}$ samples selected from each of $\mathcal{S}_{\text{rare}}$ and $\mathcal{S}_{\text{medium}}$, and $0.5N_{\mathcal{D}^{\prime}}$ samples from $\mathcal{S}_{\text{common}}$. Afterwards, we pre-train the DegDiT model on the curated dataset $\mathcal{D}^{\prime}$, enabling the model to learn robust representations from a large and diverse set of audio-text-event triplets. Subsequently, we fine-tune the model on the AudioCondition [21] training set, which provides high-quality, human-annotated event descriptions and temporal boundaries. This two-stage training procedure allows the model to benefit from both large-scale weakly labeled data and precise, expert-annotated examples, enhancing its generalization and controllability in TTA generation tasks.
@@ III-C Consensus Preference Optimization
Generative models are increasingly trained using reinforcement learning from human feedback (RLHF) [38, 39, 40] to better align model outputs with human preferences. In this paper, controllable audio generation poses unique challenges, as it requires balancing multiple reward criteria—including timestamp accuracy, event matching, and audio quality. To address this, we introduce a multi-reward reference optimization strategy designed to integrate these diverse objectives. We first pick 5k audios and their corresponding prompts from AudioTime [28], which is a manually generated dataset with accurate timestamps. Utilizing the same prompts, we generate 5k audios using TangoFlux [17] and the fine-tuned DegDiT, respectively. Leveraging these three parts of data, we extend the standard preference learning by incorporating fine-grained, multi-dimensional quality assessments named Consensus Preference Optimization (CoPO). Unlike Direct Preference Optimization (DPO) [38], which treats preferences as binary signals, CoPO leverages continuous preference intensities derived from multiple complementary reward metrics, since multiple evaluation criteria are available in controllable audio generation tasks. In detail, the reward components are computed through specialized functions: text alignment via CLAP-score between audio and text embeddings ($r_{\text{text}}$), event alignment through comparison with parsed event descriptions ($r_{\text{event}}$), temporal accuracy against ground truth timestamps ($r_{\text{temporal}}$), and audio quality combining signal-to-noise ratio, spectral balance and dynamic range metrics ($r_{\text{audio}}$). These components are aggregated through the weights $\mathbf{w}=[w_{\text{text}},w_{\text{event}},w_{\text{temporal}},w_{\text{audio}}]$ to form an overall quality score:
@@
The complete CoPO objective combines preference learning with reconstruction quality:
@@
with $\beta$ controlling sensitivity to preference differences and $\lambda$ balancing the two objectives. The training procedure alternates between computing multi-dimensional rewards, updating model parameters, and periodically refreshing the reference model through exponential moving averaging.
@@ IV-A Implementations Details
Datasets. We conduct our experiments using a diverse set of publicly available datasets to ensure comprehensive evaluation. For pre-training, we utilize AudioSet [26], a large-scale, weakly-labeled dataset comprising over 2 million 10-second YouTube video clips annotated with 527 sound event classes. We apply the Quality-Balanced Data Selection pipeline to produce a curated subset of $N_{\mathcal{D}^{\prime}}=100,000$ high-quality samples for model pre-training. For fine-tuning, we employ the AudioCondition training set [21], including 89,557 strongly-labeled audio pieces with 439 event categories, which leverage audio and timestamp annotations from AudioSetStrong [41]. During the CoPO phase, we leverage the AudioTime dataset [28], which offers manually annotated audio samples with precise temporal information for sound events. We select 5,000 samples with corresponding prompts and generate 5,000 synthetic audio samples using TangoFlux [17] and our fine-tuned DegDiT model employing the same prompts to construct paired real and synthetic samples. For evaluation, we utilize the test split of AudioCondition [21], comprising 1,110 audio samples, as well as 692 samples from the DESED [27] test set, which contains 10 sound event classes with strong temporal annotations for assessing temporal control accuracy. We also incorporate the test sets of AudioTime [28], which offers high-quality audio with explicit event borders and a diverse range of event types, suitable for subjective evaluation.
@@
We also design Audio Evaluation Instructions in Fig. 3 and conduct subjective evaluations with both Gemini 2.5 Pro LLM [47] and human raters. The generated samples on are rated by the LLM on content matching, timing accuracy, audio quality, and overall coherence. With the same four metrics, we also engage 10 domain experts in a subjective evaluation, where each expert independently rates 30 randomly selected samples from AudioTime [28] for each model.
@@
IV-B Comparison with SOTA methods
DegDiT’s performance advantage extends to subjective evaluation as well. According to the Gemini 2.5 Pro Rating, DegDiT consistently receives the highest or near-highest scores across content matching, timing accuracy, and audio quality metrics on both datasets. Its elevated Overall scores suggest that the model not only excels in specific aspects but also provides a well-rounded and satisfying listening experience. By comparison, while methods such as AudioLDM2, TangoFlux (FT), and AudioComposer may perform competitively on certain individual metrics, they do not achieve the same level of balanced performance across all evaluation criteria. These findings underscore DegDiT’s effectiveness in advancing the state of controllable audio generation, both in terms of temporal accuracy and semantic alignment.
@@ IV-B Comparison with SOTA methods
In Tab. IV, we present a comparison of the audio generation models on the AudioTime test set, evaluated across both objective and subjective metrics. The results demonstrate that DegDiT consistently outperforms all the baselines, achieving the best performance in the majority of objective metrics, including duration, frequency, ordering, and timestamp accuracy. This indicates that DegDiT is highly effective at capturing the temporal structure and event alignment present in the reference audio. Furthermore, DegDiT attains the highest scores in all subjective evaluation categories—content, temporal coherence, audio quality, and overall impression—indicating a significant advancement in perceptual audio generation quality. Notably, the improvements in both objective and subjective metrics suggest that DegDiT produces audio that is structurally faithful and delivers outputs that are more natural and convincing to human listeners.
In Fig. 4, we present representative examples of prompts containing single events, two events, multiple events, and overlapping events for three models: Tangoflux, Audiocomposer, and our proposed DegDiT. Tangoflux does not incorporate explicit timestamp control signals, which limits its ability to generate audio with accurate temporal alignment of events. In contrast, Audiocomposer relies solely on textual descriptions for control. As a result, it sometimes fails to capture all events in multi-event scenarios and struggles to handle cases where events overlap in time. By comparison, our proposed DegDiT demonstrates robust performance across all these challenging conditions, effectively generating audio that accurately reflects the specified event timing and content, even in the presence of multiple or overlapping events.
@@
IV-C Ablation Study
In this section, we present the ablation study and parameter sensitivity experiments for our proposed method. As shown in Tab. V, the progressive increases in both F1 (Event) and F1 (Clip) scores demonstrate the effectiveness of each proposed module. We initially train the diffusion transformer on the AudioCondition dataset, denoted as DiT (FT), following the fine-tuning procedure described earlier. The substantial improvement following the introduction of Dynamic Event Graphs (DEG) underscores the importance of modeling temporal dependencies and event interactions, enabling the system to more accurately capture the complex structure of audio events. Further gains achieved through the quality-balanced data selection (QBDS) pipeline suggest that careful curation of training data can effectively mitigate the adverse effects of noise and class imbalance. Finally, the additional improvements observed with Consensus Preference Optimization (CoPO) indicate that leveraging the consensus among multiple reward models leads to comprehensive and robust enhancements in model performance. These results highlight the significance of integrating temporal modeling, data quality control, and consensus-driven optimization to advance the state of audio event detection.
@@ IV-C Ablation Study
When performing inference with diffusion models, adjusting the guidance scale and the number of diffusion steps can influence output fidelity, diversity, and computational resources required. According to Tab. VII, when the guidance scale is set to zero, the model fails to produce the target audio under DEG guidance. However, as the guidance scale increases within a reasonable range, only moderate changes are observed in event-level and clip-level F1 scores, as well as in perceptual and audio-text alignment metrics. Similarly, varying the number of diffusion denoising steps does not lead to significant changes in any of the reported metrics. Across both the AudioCondition and DESED datasets, the model maintains robust performance under all tested configurations, with only slight variations. These results indicate that the model’s effectiveness in generating high-quality, semantically accurate audio remains stable, even when inference-time parameters are adjusted within a reasonable range.
@@ V Conclusion and future work
In this paper, we propose DegDiT, a dynamic event graph guided diffusion transformer framework for controllable text-to-audio generation. By encoding multi-event prompts as structured event graphs and modeling both semantic and temporal relationships, DegDiT enables fine-grained control over generated audio content. The introduction of a quality-balanced data selection pipeline and consensus preference optimization further enhances data diversity and multi-reward alignment. Experimental results on several benchmarks demonstrate that DegDiT achieves state-of-the-art performance in terms of flexibility, temporal precision, and usability, establishing a new paradigm for controllable audio synthesis.
Despite its strong performance, DegDiT occasionally generates redundant audio segments when faced with scenarios involving rare or uncommon events. This issue is likely attributable to the lack of event categories within the training dataset. Due to this, our future work will focus on constructing a large-scale dataset with precise timestamp annotations covering a diverse array of events. We anticipate that such a dataset will not only enhance the performance of controllable text-to-audio generation, especially for rare events, but also facilitate more robust and reliable audio synthesis across a broader range of real-world scenarios.

### v2

@@ III-B Training Strategy and Dataset Construction
$\mathbf{E}_{\text{cond}}=\textrm{Concat}(\mathbf{E}_{\text{text}},\mathbf{H},\mathbf{E}_{\text{dur}}).$ (9)
@@
$\mathbf{z}_{t}=(1-t)\mathbf{x}_{1}+t\mathbf{x}_{0},$ (10)
@@
$\mathbf{u}_{t}=\frac{d\mathbf{z}_{t}}{dt}=\mathbf{x}_{0}-\mathbf{x}_{1}.$ (11)
Enabling the model to learn a time-dependent vector field that transports samples from the noise distribution to the target audio distribution, we train the model $\epsilon_{\theta}(\mathbf{z}_{t},t,\mathbf{E}_{\text{cond}})$ to predict this velocity using the flow matching loss:
$\mathcal{L}_{\text{FM}}=\mathbb{E}_{\mathbf{x}_{1},\mathbf{x}_{0},t}\left[\|\epsilon_{\theta}(\mathbf{z}_{t},t,\mathbf{E}_{\text{cond}})-\mathbf{u}_{t}\|_{2}^{2}\right].$ (12)
@@
Let $N_{\mathcal{D}^{\prime}}$ denotes the predefined sample numbers of $\mathcal{D}^{\prime}$, with $0.25N_{\mathcal{D}^{\prime}}$ samples selected from each of $\mathcal{S}_{\text{rare}}$ and $\mathcal{S}_{\text{medium}}$, and $0.5N_{\mathcal{D}^{\prime}}$ samples from $\mathcal{S}_{\text{common}}$. Afterwards, we pre-train the DegDiT model on the curated dataset $\mathcal{D}^{\prime}$, enabling the model to learn robust representations from a large and diverse set of audio-text-event triplets. Subsequently, we fine-tune the model on the AudioCondition [18] training set, which provides high-quality event descriptions and temporal boundaries. This two-stage training procedure allows the model to benefit from both large-scale weakly labeled data and precise, expert-annotated audio examples.
@@ III-C Consensus Preference Optimization
Generative models are increasingly trained using reinforcement learning from human feedback (RLHF) [39, 2, 31] to better align model outputs with human preferences. In this paper, controllable audio generation poses unique challenges, as it requires balancing multiple reward criteria—including timestamp accuracy, event matching, and audio quality. To address this, we introduce a multi-reward reference optimization strategy designed to integrate these diverse objectives. We first pick 5k audios and their corresponding prompts from AudioTime [50], which is a manually generated dataset with accurate timestamps. Utilizing the same prompts, we generate 5k audios using TangoFlux [23] and the fine-tuned DegDiT, respectively. With these three parts of data, we extend the standard preference learning by incorporating fine-grained, multi-dimensional quality assessments named Consensus Preference Optimization (CoPO). Unlike Direct Preference Optimization (DPO) [39], which treats preferences as binary signals, CoPO leverages continuous preference intensities derived from multiple reward metrics, since multiple evaluation criteria are available in controllable audio generation tasks. In detail, the reward components are computed through specialized functions: text alignment via CLAP-score between audio and text embeddings ($r_{\text{text}}$), event alignment through comparison with parsed event descriptions ($r_{\text{event}}$), temporal accuracy against ground truth timestamps ($r_{\text{temporal}}$), and audio quality combining signal-to-noise ratio, spectral balance and dynamic range metrics ($r_{\text{audio}}$). These components are aggregated through the weights $\mathbf{w}=[w_{\text{text}},w_{\text{event}},w_{\text{temporal}},w_{\text{audio}}]$ to form an overall quality score:
@@
The complete CoPO objective combines preference learning loss with a flow matching objective:
@@
with $\beta$ controlling sensitivity to preference differences and a trade-off parameter $\lambda$. The training procedure iteratively alternates between two steps: computing multi-dimensional rewards and updating the model parameters. In parallel, the reference model is periodically refreshed through exponential moving averaging.
@@ IV-A Implementations Details
Datasets. We conduct our experiments using a diverse set of publicly available datasets to ensure comprehensive evaluation. For pre-training, we utilize AudioSet [15], a large-scale, weakly-labeled dataset comprising over 2 million 10-second YouTube video clips annotated with 527 sound event categories. We apply the Quality-Balanced Data Selection pipeline to produce a curated subset of $N_{\mathcal{D}^{\prime}}=100,000$ high-quality samples for model pre-training. For fine-tuning, we employ the AudioCondition training set [18], including 89,557 strongly-labeled audio pieces with 439 event categories, which leverage audio and timestamp annotations from AudioSetStrong [21]. During the CoPO phase, we leverage the AudioTime dataset [50], which offers manually annotated audio samples with precise temporal information for sound events. We select 5,000 samples with corresponding prompts and generate 5,000 synthetic audio samples using TangoFlux [23] and our fine-tuned DegDiT model employing the same prompts to construct paired real and synthetic samples. For evaluation, we utilize the test split of AudioCondition [18], comprising 1,110 audio samples, as well as 692 samples from the DESED [44] test set, which contains 10 sound event classes with strong temporal annotations for assessing temporal control accuracy. We also incorporate the test sets of AudioTime [50], which offers high-quality audio with explicit event borders and a diverse range of event types, suitable for subjective evaluation.
@@
Since experiments in WoW-Bench [27] show that MLLMs perform comparably to human raters in judging audio categories and duration, we design Audio Evaluation Instructions in Fig. 3 and conduct subjective evaluations with both Gemini 2.5 Pro MLLM [9] and human raters. The generated samples are rated by the MLLM on content matching, timing accuracy, audio quality, and overall coherence. With the same four metrics, we also engage 17 domain experts in a subjective evaluation, where each expert independently rates 30 randomly selected samples from AudioTime [50] for each model. The Mean Opinion Score (MOS) along with the 95% confidence intervals are computed for the results.
@@
During inference, we employ Classifier-Free Guidance (CFG) to enhance the semantic and temporal fidelity of the generated audio. The conditional model $\epsilon_{\theta}(\mathbf{z}_{t},t,\mathbf{E}_{\text{cond}})$ and unconditional model $\epsilon_{\theta}(\mathbf{z}_{t},t,\emptyset)$ are combined to compute the guided velocity prediction:
$\begin{split}\tilde{\epsilon}_{\theta}(\mathbf{z}_{t},t,\mathbf{E}_{\text{cond}})=&\epsilon_{\theta}(\mathbf{z}_{t},t,\emptyset)\\ &+\gamma\cdot\left(\epsilon_{\theta}(\mathbf{z}_{t},t,\mathbf{E}_{\text{cond}})-\epsilon_{\theta}(\mathbf{z}_{t},t,\emptyset)\right),\end{split}$ (18)
where $\gamma$ denotes the guidance scale. We find $\gamma=4.0$ to be optimal in our experiments. The inference process follows the rectified flow ODE, which is solved numerically using the guided velocity $\tilde{\epsilon}_{\theta}$ to produce audio samples.
@@
IV-B Comparison with SOTA methods
DegDiT’s performance also extends to evaluations by both large audio-language models and human evaluators. According to the Gemini 2.5 Pro Rating, DegDiT consistently receives the highest or near-highest scores across content matching, timing accuracy, and audio quality metrics on both datasets. Its elevated Overall scores suggest that the model not only excels in specific aspects but also provides a satisfying listening experience. By comparison, while methods such as AudioLDM2, TangoFlux (FT), and AudioComposer may perform competitively on certain individual metrics, they do not achieve the same level of balanced performance across all evaluation criteria. These findings underscore DegDiT’s effectiveness in advancing the state of controllable audio generation, both in terms of temporal accuracy and semantic alignment.
@@ IV-B Comparison with SOTA methods
In Tab. IV, we present a comparison of the audio generation models on the AudioTime test set, evaluated across both objective and subjective metrics. The results demonstrate that DegDiT consistently achieves the best performance in the majority of objective metrics, including duration, frequency, ordering, and timestamp accuracy. This indicates that DegDiT is highly effective at capturing the temporal structure and event alignment present in the reference audio. Furthermore, DegDiT attains the highest scores in all subjective evaluation categories—content, temporal coherence, audio quality, and overall impression. Notably, the improvements in both objective and subjective metrics suggest that DegDiT produces audio that is natural and convincing to human listeners.
In Fig. 4, we present representative examples of prompts containing single events, two events, multiple events, and overlapping events for three models: Tangoflux, Audiocomposer, and our proposed DegDiT. Tangoflux does not incorporate explicit timestamp control signals, which limits its ability to generate audio with accurate temporal alignment of events. In contrast, Audiocomposer relies solely on textual descriptions for control. As a result, it sometimes fails to capture all events in multi-event scenarios and struggles to handle cases where events overlap in time. By comparison, our proposed DegDiT demonstrates robust performance across all these challenging conditions and generates audio that accurately reflects the specified event timing and content.
IV-C Ablation Study
@@
In this section, we present the ablation study and parameter sensitivity experiments for our proposed method. As shown in Tab. V, the progressive increases in both F1 (Event) and F1 (Clip) scores demonstrate the effectiveness of each proposed module. We initially train the diffusion transformer on the AudioCondition dataset, denoted as DiT (FT), following the fine-tuning procedure described earlier. The substantial improvement following the introduction of Dynamic Event Graphs (DEG) underscores the importance of modeling temporal dependencies and event interactions, enabling the system to more accurately capture the complex structure of audio events. Further gains achieved through the Quality-Balanced Data Selection (QBDS) pipeline suggest that careful curation of training data can mitigate the adverse effects of noise and class imbalance. Finally, the additional improvements observed with Consensus Preference Optimization (CoPO) indicate that leveraging the consensus among multiple reward models leads to comprehensive enhancements in model performance. Consequently, our findings clarify that the DEG embedding cannot function as a standalone control signal, as experiments show that training without a text prompt leads to non-convergence.
@@
When performing inference with diffusion models, adjusting the guidance scale and the number of diffusion steps can influence output fidelity, diversity, and computational resources required. According to Tab. VII, when the guidance scale $\gamma$ is set to zero, the model fails to produce the target audio under DEG guidance. However, as $\gamma$ increases within a reasonable range, only moderate changes are observed in event-level and clip-level F1 scores, as well as in perceptual and audio-text alignment metrics. Similarly, varying the number of diffusion denoising steps does not lead to significant changes in any of the reported metrics. Across both the AudioCondition and DESED datasets, the model maintains robust performance under all tested configurations, with only slight variations. These results indicate that the model’s effectiveness in generating high-quality, semantically accurate audio remains stable, even when inference-time parameters are adjusted within a reasonable range.
TABLE VIII: Comparison of Model Complexity and Inference Efficiency.
Method Params (M) GPU Mem (GB) Time (s)
AudioLDM2 346 2.46 17.54
Tango2 866 5.39 17.35
TangoFlux* 515 5.11 2.76
AudioComposer 272 10.22 1.56
DegDiT (Ours) 568 5.31 2.87
IV-D Computational Requirements
[…]
@@ V Conclusion and future work
In conclusion, we propose DegDiT, a dynamic event graph guided diffusion transformer framework for controllable text-to-audio generation. By encoding multi-event prompts as structured event graphs and modeling their temporal relationships, DegDiT enables fine-grained control over generated audio content. The introduction of a quality-balanced data selection pipeline and consensus preference optimization further enhances data diversity and multi-reward alignment. Experimental results on several benchmarks demonstrate that DegDiT achieves state-of-the-art performance in terms of temporal precision and usability.
Despite its strong performance, DegDiT occasionally generates redundant audio segments when faced with scenarios involving the minority event classes [7]. To understand the causes, we conducted an analysis based on Stable Flow [1]. It reveals that while generating these minority events, attention leakage occurs in later steps and deeper transformer layers, leading to redundant output. Due to this, our future work will focus on constructing a large-scale dataset with precise timestamp annotations covering a diverse array of events. We anticipate that such a dataset will not only enhance the performance of controllable text-to-audio generation, especially for rare events, but also facilitate more robust and reliable audio synthesis across a broader range of real-world scenarios.